OpenStudy (anonymous):

write e^ln(8.1) with out using "e" or ln

OpenStudy (lalaly):

\[\huge{e^{\ln(8.1)}=8.1}\]

OpenStudy (lalaly):

\[\huge{e^{\ln(x)}=x}\]

OpenStudy (anonymous):

what ln(x) means is that you're for a number y where e^y = x. They are "opposite" operations in a way and, because of this, it's the same thing as having 8.1+2-2 or 8.1*2/2.

OpenStudy (anonymous):

Think about what \(\ln8.1\) means. It means that \(e^x=8.1\), and it returns the value \(x\). So, if you're taking \(e^{\ln{8.1}}\) , then you're just taking it to that same number, so it has to equal 8.1